carson035:

solve on the interval 0 2pi) 2secx+5=1

carson035:

Shadow:

Hello

Shadow:

You start by isolating x, as you would in algebra

Shadow:

\[2\sec x + 5 = 1\] \[2 \sec x = -4\] \[\sec x = -2\] Since: \[\sec = \frac{ 1 }{ \cos }\] We get: \[\frac{ 1 }{ \cos x} = -2\]

Shadow:

Do you know what to do from there?

Shadow:

\[1 = -2 \cos x\] \[\cos x = -\frac{ 1 }{ 2 }\]

Shadow:

@carson035

Shadow:

Do you know what you would do from here?

Shadow:

Cos of what angle = -1/2 Cosine refers to the x coordinate of an angle, and since this is negative, you want to look to the left side of the unit circle (quadrants 2 and 3)

Shadow:

Once you do that, the answer is pretty simple to find

Shadow:

\[x = \frac{ 2 \pi }{ 3 }, \frac{ 4 \pi }{ 3 }\]

Shadow:



Shadow:

These guys

Shadow:

@carson035